# Debiased Inference for Dynamic Nonlinear Panels with Multi-dimensional Heterogeneities

### v3

@@ 2.1 DN-HP Model
Consider a probit model
$Y_{it}=\mathds{1}\left\{\pi_{it}\left(\theta,\alpha_{i},\gamma_{t}\right)+\varepsilon_{it}>0\right\}$
where the linear index $\pi_{it}(\cdot)$ varies across both individuals and time periods. The error term $\varepsilon_{it}$ follows a standard normal distribution denoted by $\Phi$. Without of loss of generality, let $X_{it}\in\mathbb{R},$ and the parameter vectors $\theta=\left(\theta_{1},\theta_{2}\right)^{\prime},$ $\alpha_{i}=\left(\alpha_{1,i},\alpha_{2,i}\right)^{\prime},$ and $\gamma_{t}=\left(\gamma_{1,t},\gamma_{2,t}\right)^{\prime}.$ We accommodate the linear index of the form:
$\pi_{it}\left(\theta,\alpha_{i},\gamma_{t}\right)=\left(\theta_{1}+\alpha_{1,i}+\gamma_{1,t}\right)+X_{it}\left(\theta_{2}+\alpha_{2,i}+\gamma_{2,t}\right).$ (2.1)
In the special case where $\alpha_{i}=\alpha_{1,i}$ and $\gamma_{t}=\gamma_{1,t},$ (2.1) simplifies to
$\pi_{it}\left(\theta,\alpha_{i},\gamma_{t}\right)=\left(\theta_{1}+\alpha_{1,i}+\gamma_{1,t}\right)+X_{it}\theta_{2},$
which corresponds to the traditional two-way fixed effects model. Assume that $\varepsilon_{it}$ is independent of $X_{it}.$ The likelihood for individual $i$ at time $t$ is given by
$l_{i,t}\left(\theta,\phi\right):=Y_{it}\log\Phi\left(\pi_{it}\left(\theta,\alpha_{i},\gamma_{t}\right)\right)+\left(1-Y_{it}\right)\log\Phi\left(-\pi_{it}\left(\theta,\alpha_{i},\gamma_{t}\right)\right).$
Due to the additive form of the parameters, the structural parameters, $\theta,$ is not identified in the likelihood. To resolve this, we reparameterize the likelihood by setting $\alpha_{N}=-\sum_{i=1}^{N-1}\alpha_{i}$ and $\gamma_{T}=-\sum_{t=1}^{T-1}\gamma_{t},$ thereby imposing the constraints $\sum_{i=1}^{N}\alpha_{i}=0$ and $\sum_{t=1}^{T}\gamma_{t}=0.$ Let $\psi:=\left(\alpha_{1}^{\prime},\ldots,\alpha_{N-1}^{\prime},\gamma_{1}^{\prime},\ldots,\gamma_{T-1}^{\prime}\right)^{\prime}.$ The reparameterized likelihood is then $l\left(\theta,D^{\prime}\psi\right)=:l^{*}\left(\theta,\psi\right)$ for $D:=\operatorname*{diag}\{D_{1},D_{2}\}$ where $D_{1}:=(\mathbb{I}_{N-1},-\iota_{N-1})\otimes\mathbb{I}_{K}$ and $D_{2}:=(\mathbb{I}_{T-1},-\iota_{T-1})\otimes\mathbb{I}_{K}.$
@@ 2.2 Incidental Parameter Problem
In this section, we illustrate the incidental parameter problem (IPP) that arises in our heterogeneous parameter model. For a given $\theta,$ define
$\psi(\theta):=\arg\max_{\psi\in\Lambda_{\theta}}\mathbb{E}[l^{*}(\theta,\psi)],\quad\widehat{\psi}(\theta):=\arg\max_{\psi\in\Lambda_{\theta}}l^{*}(\theta,\psi)$ (2.2)
where $\Lambda_{\theta}$ is a compact subset of $\mathbb{R}^{(N+T-2)K}$ for fixed $N$ and $T$ (Assumption 1) and $\mathbb{E}$ denotes expectations with respect to the true conditional density. Specifically, $\mathbb{E}$ represents the conditional expectation given the true values of fixed effects and the covariates. For notational simplicity, we use $\mathbb{E}$ when no confusion arises. Let the score and Hessian be
$s\left(\theta,\psi\right):=\partial_{\psi}l^{*}\left(\theta,\psi\right),\quad H\left(\theta,\psi\right):=\partial_{\psi\psi^{\prime}}l^{*}\left(\theta,\psi\right),$
where $\partial_{\psi}$ and $\partial_{\psi\psi^{\prime}}$ represent the first and second partial derivatives with respect to $\psi$. By (2.2), we know that $s(\theta,\widehat{\psi}\left(\theta\right))=0.$ Given $\theta,$ consider a Taylor expansion of $s(\theta,\widehat{\psi}\left(\theta\right))$ around $\psi(\theta)$:
$0=s\left(\theta,\psi\left(\theta\right)\right)+H\left(\theta,\psi(\theta)\right)\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}+r^{s}(\theta,a),$ (2.3)
where
$r^{s}\left(\theta,a\right)$ $:=\big{(}r_{\psi_{1}}^{s}\left(\theta,a_{1}\right),\ldots,r_{\psi_{(N+T-2)K}}^{s}(\theta,a_{(N+T-2)K})\big{)}^{\prime},$

### v4

@@ 2.1 Models and Estimations
Consider a probit model with data generating process (DGP)
$Y_{it}|X_{i1},\ldots,X_{it},\phi^{0}=\mathds{1}\{\pi(X_{it},\theta_{0},\alpha_{i}^{0},\gamma_{t}^{0})+\varepsilon_{it}>0\},$ (2.2)
where $\varepsilon_{it}$ is a standard normal random variable independent of $X_{it}$, $\theta_{0}$, $\alpha_{i}^{0}$, and $\gamma_{t}^{0}$. For arguments $\theta$, $\alpha_{i}$, and $\gamma_{t}$ of dimension $K\times 1$, we allow the linear index [^3] We use a model with a linear index solely for the purpose of illustration. Our method applies to general models. $\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ to take the following general form:
$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})=\sum_{k=1}^{K}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it},$
where $X_{k,it}$, $\theta_{k}$, $\alpha_{k,i}$, and $\gamma_{k,t}$ are the $k$-th components of $X_{it}$, $\theta$, $\alpha_{i}$, and $\gamma_{t}$, respectively, $k=1,\ldots,K$. In addition, when $X_{1,it}=1$, the DGP in (2.2) consists of the traditional individual- and time-effects in the intercept. When $X_{k,it}=Y_{it-1}$ for some $k$, the DGP is dynamic because each $Y_{it}$ depends on its past value.
Moreover, we also allow a subset of the coefficients to be homogeneous, thereby accommodating both heterogeneous and homogeneous parameters. For instance, when it is known that $\alpha_{1,i}^{0}=\gamma_{1,t}^{0}=0$, the model can be estimated under constraints $\alpha_{1,i}=\gamma_{1,t}=0$ for all $(i,t)$. In this case, the linear index is effectively
$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})=\theta_{1}X_{1,it}+\sum_{k=2}^{K}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it},$
where $X_{1,it}$ is associated with a “homogeneous parameter” $\theta_{1}$ while all other $X_{k,it}$ ($k=2,\ldots,K$) are associated with heterogeneous parameters $(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})$. In a similar fashion, our approach also applies to, e.g.,
$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})=(\alpha_{1,i}+\gamma_{1,t})X_{1,it}+\sum_{k=2}^{K}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it},$
with the constraint $\theta_{1}=0$. Other variants of $\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ include:
$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ $=(\alpha_{1,i}+\gamma_{1,t})+\theta_{2}Y_{it-1}+\sum_{k=3}^{K}\theta_{k}X_{k,it},$ (2.3)
$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ $=(\theta_{1}+\alpha_{1,i}+\gamma_{1,t})+\sum_{k=2}^{K}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it},$ (2.4)
$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ $=(\theta_{1}+\alpha_{1,i}+\gamma_{1,t})+(\theta_{2}+\alpha_{2,i}+\gamma_{2,t})Y_{it-1}$
$\quad+\sum_{k=3}^{K_{1}}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it}+\sum_{k=K_{1}+1}^{K}\theta_{k}X_{k,it}$ (2.5)
for some $3\leq K_{1}<K$. To our knowledge, except (2.3) considered by Fernández-Val and Weidner (2016), none of the above specifications are fully addressed under nonlinear models with predetermined regressors. The specification in (2.4) is the same to Keane and Neal (2020) and Lu and Su (2023), except that they consider only the linear regression model. Jochmans and Otsu (2019) provide a bias correction technique for general nonlinear models with (2.4) only for strictly exogenous $X_{k,it}$. Currently, no methods are available for nonlinear models with (2.5).
We are interested in the estimation of and the inference about $\theta_{0}$ in (2.1), under the presence of the nuisance parameter $\phi^{0}$. Our approach is likelihood-based. Denote $\phi:=(\alpha_{1}^{\prime},\ldots,\alpha_{N}^{\prime},\gamma_{1}^{\prime},\ldots,\gamma_{T}^{\prime})^{\prime}$ and define the log-likelihood function (“likelihood” hereafter) as
$\ell(\theta,\phi):=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}l_{i,t}(\theta,\phi),\qquad l_{i,t}(\theta,\phi):=\log f(Y_{it}|X_{it},\theta+\alpha_{i}+\gamma_{t}).$ (2.6)
Due to the additive structure $\theta+\alpha_{i}+\gamma_{t}$, $\theta$ is not identified because the likelihood function is invariant to transformations
$\alpha_{i}\mapsto\alpha_{i}+c_{1},\qquad\gamma_{t}\mapsto\gamma_{t}+c_{2},\qquad\theta\mapsto\theta-c_{1}-c_{2}$ (2.7)
for all pairs of constant $(c_{1},c_{2})$ and all $(i,t)$. To resolve this, we follow Lu and Su (2023) to reparameterize the likelihood by setting $\alpha_{N}=-\sum_{i=1}^{N-1}\alpha_{i}$ and $\gamma_{T}=-\sum_{t=1}^{T-1}\gamma_{t}$, thereby imposing the normalization $\sum_{i=1}^{N}\alpha_{i}=0$ and $\sum_{t=1}^{T}\gamma_{t}=0$. Any “average effect” is absorbed by $\theta$. Let $\psi:=\left(\alpha_{1}^{\prime},\ldots,\alpha_{N-1}^{\prime},\gamma_{1}^{\prime},\ldots,\gamma_{T-1}^{\prime}\right)^{\prime}$, the reparameterized likelihood can be constructed from the original $\ell\left(\theta,\phi\right)$ as
$\ell\left(\theta,D^{\prime}\psi\right)=:l(\theta,\psi),\qquad D:=\operatorname*{diag}\{D_{1},D_{2}\},$
where $D_{1}:=(\mathbb{I}_{N-1},-\iota_{N-1})\otimes\mathbb{I}_{K}$ and $D_{2}:=(\mathbb{I}_{T-1},-\iota_{T-1})\otimes\mathbb{I}_{K}$. This rules out the indeterminancy in (2.7) without explicitly using the Lagrangian. We use the following example to illustrate the reparameterization.
Example 2 (reparameterization)
For some scalar structural parameter $\mu_{0}$ of interests and $K=1$, consider $Y_{it}=\mu_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}+\varepsilon_{it}$ where $\varepsilon_{it}$ is standard normal independent across both $i$ and $t$, and independent of $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$. We restrict the variance of $\varepsilon_{it}$ to be $1$ because we use this example only to demonstrate the identification of $\mu_{0}$. In Example 3 below, we indeed allow $\varepsilon_{it}$ to have an unknown variance.
Conditioning on $\phi^{0}$, the original likelihood function is
$\ell(\mu,\phi)$ $=-\frac{1}{2}\log(2\pi)-\frac{1}{NT}\sum_{i=1}^{N-1}\sum_{t=1}^{T-1}\frac{1}{2}(Y_{it}-\mu-\alpha_{i}-\gamma_{t})^{2}-\frac{1}{NT}\sum_{t=1}^{T-1}\frac{1}{2}(Y_{Nt}-\mu-\alpha_{N}-\gamma_{t})^{2}$
$\quad-\frac{1}{NT}\sum_{i=1}^{N-1}\frac{1}{2}(Y_{iT}-\mu-\alpha_{i}-\gamma_{T})^{2}-\frac{1}{NT}\frac{1}{2}(Y_{NT}-\mu-\alpha_{N}-\gamma_{T})^{2}.$
Setting $\alpha_{N}=-\sum_{i=1}^{N-1}\alpha_{i}$ and $\gamma_{T}=-\sum_{t=1}^{T-1}\gamma_{t}$, the reparameterized likelihood is
$l(\mu,\psi)$ $=-\frac{1}{2}\log(2\pi)-\frac{1}{NT}\sum_{i=1}^{N-1}\sum_{t=1}^{T-1}\frac{1}{2}(Y_{it}-\mu-\alpha_{i}-\gamma_{t})^{2}-\frac{1}{NT}\sum_{t=1}^{T-1}\frac{1}{2}\left(Y_{Nt}-\mu+\sum_{i=1}^{N-1}\alpha_{i}-\gamma_{t}\right)^{2}$
$\quad-\frac{1}{NT}\sum_{i=1}^{N-1}\frac{1}{2}\left(Y_{iT}-\mu-\alpha_{i}+\sum_{t=1}^{T-1}\gamma_{t}\right)^{2}-\frac{1}{NT}\frac{1}{2}\left(Y_{NT}-\mu+\sum_{i=1}^{N-1}\alpha_{i}+\sum_{t=1}^{T-1}\gamma_{t}\right)^{2}.$
Calculating the first derivative of $l(\mu,\psi)$ w.r.t. $\alpha_{i}$, we have
$\partial_{\alpha_{i}}l(\mu,\psi)=\frac{1}{NT}\sum_{t=1}^{T}(Y_{it}-\mu-\alpha_{i})-\frac{1}{NT}\sum_{t=1}^{T}\left(Y_{Nt}-\mu+\sum_{i=1}^{N-1}\alpha_{i}\right)$
for every $i=1,\ldots,N-1$. Equating $\partial_{\alpha_{i}}l(\mu,\psi)=0$, we obtain
$\alpha_{i}=\frac{1}{T}\sum_{t=1}^{T}Y_{it}-\frac{1}{T}\sum_{t=1}^{T}Y_{Nt}-\sum_{i=1}^{N-1}\alpha_{i}.$ (2.8)
Notice that (2.8) is independent of $\mu$ and all $\gamma_{t}$. Now, sum (2.8) over $i=1,\ldots,N-1$, we obtain
$\sum_{i=1}^{N-1}\alpha_{i}=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it}-\frac{1}{T}\sum_{t=1}^{T}Y_{Nt}.$ (2.9)
Combining (2.8) and (2.9), it follows that
$\widehat{\alpha}_{i}=\frac{1}{T}\sum_{t=1}^{T}Y_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it},\qquad i=1,\ldots,N-1.$
It can be calculated in the same way that
$\widehat{\gamma}_{t}=\frac{1}{N}\sum_{i=1}^{N}Y_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it},\qquad t=1,\ldots,T-1.$
Subsequently,
$\widehat{\mu}=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it}.$ (2.10)
In this simple example, $\widehat{\alpha}_{i}$ and $\widehat{\gamma}_{t}$ do not depend on $\mu$. This is generally not the case for complex models.
In addition, maximizing $l(\mu,\psi)$ is identical to maximizing the original likelihood $\ell\left(\mu,\phi\right)$ under constraints $\sum_{i=1}^{N}\alpha_{i}=0$ and $\sum_{t=1}^{T}\gamma_{t}=0$. Specifically, it can be verified that the Lagrangian $\ell\left(\mu,\phi\right)+\lambda^{\alpha}\sum_{i=1}^{N}\alpha_{i}+\lambda^{\gamma}\sum_{t=1}^{T}\gamma_{t}$, where $\lambda^{\alpha}$ and $\lambda^{\gamma}$ are the two Lagrange multipliers, delivers upon maximization the same $\widehat{\alpha}_{i}$ and $\widehat{\gamma}_{t}$ for $i=1,\ldots,N-1$ and $t=1,\ldots,T$, respectively. For $\widehat{\alpha}_{N}$ and $\widehat{\gamma}_{T}$, the Lagrangian produces
$\widehat{\alpha}_{N}=\frac{1}{T}\sum_{t=1}^{T}Y_{Nt}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it},\qquad\widehat{\gamma}_{T}=\frac{1}{N}\sum_{i=1}^{N}Y_{iT}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it}.$
Indeed, $\widehat{\alpha}_{N}=-\sum_{i=1}^{N-1}\widehat{\alpha}_{i}$ as in Equation (2.9), and it can be verified that $\widehat{\gamma}_{T}=-\sum_{t=1}^{T-1}\widehat{\gamma}_{t}$ as well.
@@ 2.2 Incidental Parameter Problem
The dynamic nonlinear heterogeneous parameter, DN-HP, models generally suffer from the incidental parameter problem (IPP). We now discuss this problem. We use $\mathbb{E}$ to denote the expectation w.r.t.
$\prod_{i=1}^{N}\prod_{t=1}^{T}f(y_{it}|X_{it},\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}),$
which is the true density evaluated at $\theta_{0}$, given: i) initial values of the predetermined regressors, ii) all strictly exogenous regressors, and iii) the unobserved effect $\phi^{0}$.
Remark 1 (expectation)
Our definition of $\mathbb{E}$ depends on the model specification and is the same to Arellano and Bonhomme (2009) and Fernández-Val and Weidner (2016). Consider the frequently used autoregressive model with $X_{it}=(Y_{it-1},Z_{it}^{\prime})^{\prime}$ for some strictly exogenous regressor $Z_{it}$. It is common that, for each $i=1,\ldots,N$, the density of $Y_{i1}$ is missing and (2.6) becomes a conditional likelihood as in Chapter 5 of Hamilton (1994). In this case, the expectation is taken w.r.t. $f_{Y_{i2},\ldots,Y_{iT}}\left(\cdot|Y_{i1},Z_{i1},\ldots,Z_{iT},\phi^{0};\theta_{0}\right)$, which is the density of only $Y_{i2},\ldots,Y_{iT}$, conditional on the initial value $Y_{i1}$, strictly exogenous regressors $Z_{i1},\ldots,Z_{iT}$, and the unobserved effects $\phi^{0}$. However, when the density of $Y_{i1}$ is indeed specified, so that (2.6) is an exact likelihood, the expectation is taken w.r.t. $f_{Y_{i1},\ldots,Y_{iT}}\left(\cdot|Z_{i1},\ldots,Z_{iT},\phi^{0};\theta_{0}\right)$, which is the joint density of all $Y_{i1},\ldots,Y_{iT}$, conditioning on the strictly exogenous regressors $Z_{i1},\ldots,Z_{iT}$ and the unobserved effects $\phi^{0}$ only.
Following the profiled likelihood framework of Pace and Salvan (2006), we continue our discussion by defining, for a given $\theta$ and each pair of $(N,T)$,
$\widehat{\psi}(\theta):=\arg\max_{\psi}l(\theta,\psi),\qquad\psi(\theta):=\arg\max_{\psi}\mathbb{E}[l(\theta,\psi)].$ (2.11)
Here $\psi(\theta)$ is referred to as the pseudo-true value and $l(\theta,\psi(\theta))$ is not subject to the IPP. See, e.g., Hahn and Newey (2004) and Arellano and Hahn (2016). In practice, however, $\psi(\theta)$ is generally infeasible and the plug-in version $l(\theta,\widehat{\psi}(\theta))$ is used instead. Specifically, the MLE of $\theta_{0}$, $\widehat{\theta}:=\arg\max_{\theta}\ell(\theta,\widehat{\psi}(\theta))$. The LR and LM test statistics (for hypothesis about $\theta_{0}$) are also constructed using $l(\theta,\widehat{\psi}(\theta))$. A fundamental problem is that $l(\theta,\widehat{\psi}(\theta))$ contains a non-negligible asymptotic bias of order $O(1/\sqrt{NT})$, due to the estimation errors in $\widehat{\psi}(\theta)$. This is referred to as the IPP in the context of the DN-HP model. The consequence is that, $\widehat{\theta}$ and many likelihood-based test statistics deviate from their standard asymptotic distribution, leading to incorrect inferences.
To demonstrate the non-negligible bias in $l(\theta,\widehat{\psi}(\theta))$, consider a Taylor expansion of $l(\theta,\widehat{\psi}(\theta))$ around $\psi(\theta)$:
$l(\theta,\widehat{\psi}(\theta))=l(\theta,\psi(\theta))+s^{\prime}(\theta)[\widehat{\psi}(\theta)-\psi(\theta)]+\frac{1}{2}[\widehat{\psi}(\theta)-\psi(\theta)]^{\prime}H(\theta)[\widehat{\psi}(\theta)-\psi(\theta)]+r^{l}(\theta),$ (2.12)
where $\widehat{\psi}(\theta)-\psi(\theta)$ represents the estimation errors of the fixed effects,
$s(\theta)$ $:=s(\theta,\psi(\theta)),\qquad s(\theta,\psi):=\partial_{\psi}l(\theta,\psi),$
$H\left(\theta\right)$ $:=H(\theta,\psi(\theta)),\qquad H(\theta,\psi):=\partial_{\psi\psi^{\prime}}l(\theta,\psi)$
are the score and Hessian. The remainder term $r^{l}(\theta)$ is characterized by Equation (A.17) in Appendix A. Here the estimation errors $\widehat{\psi}(\theta)-\psi(\theta)$ can be derived by another Taylor expansion. Seeing $s(\theta,\widehat{\psi}(\theta))=0$ as $\widehat{\psi}(\theta)$ is the maximizer, a Taylor expansion of $s(\theta,\widehat{\psi}(\theta))$ around $\psi(\theta)$ gives
00 0 $=s\left(\theta\right)+H\left(\theta\right)[\widehat{\psi}(\theta)-\psi(\theta)]+r^{s}(\theta),$
$\widehat{\psi}(\theta)-\psi(\theta)$ $=-[H\left(\theta\right)]^{-1}s\left(\theta\right)-[H\left(\theta\right)]^{-1}r^{s}(\theta),$ (2.13)
where $r^{s}(\theta)$ is the remainder term of the expansion, which is given by Equation (A.6) in Appendix A. Combining (2.13) and (2.12), and taking expectation, we obtain
$\sqrt{NT}\mathbb{E}[l(\theta,\widehat{\psi}(\theta))-l(\theta,\psi(\theta))]=-\sqrt{NT}B(\theta)+\sqrt{NT}\mathbb{E}r(\theta),\qquad B(\theta):=\frac{1}{2}\mathbb{E\{}s^{\prime}(\theta)[H(\theta)]^{-1}s(\theta)\},$ (2.14)
where $r(\theta)$ is a remainder term depending on $r^{l}(\theta)$ and $r^{s}(\theta)$, given by Equation (A.16) in Appendix A. By Lemma A.3 in Appendix A, we show that $\mathbb{E}r(\theta)$ is of order $o(1/\sqrt{NT})$, uniformly in $\theta$, as $N,T\rightarrow\infty$ with $N/T\rightarrow\kappa$ for some $0<\kappa<\infty$. This implies that $\sqrt{NT}\mathbb{E}r(\theta)=o(1)$ and is negligible. $B(\theta)$ is the IPP bias arising from the estimation errors of $\widehat{\psi}(\theta)$. As opposite to $\mathbb{E}r(\theta)$, the bias $B(\theta)$ is not negligible, in the sense that $\sqrt{NT}\mathbb{E}B(\theta)\not\rightarrow 0$ under the above asymptotic sequences $\left\{N,T\right\}$ (see Theorem 1). Consequently, $\widehat{\theta}$ and many likelihood-based test statistics inherit the bias from the likelihood, leading to invalid inferences. We consider the asymptotic sequences $N,T\rightarrow\infty$ with $N/T\rightarrow\kappa$ for some $0<\kappa<\infty$, because $\sqrt{NT}B(\theta)$ explodes under $N/T\rightarrow 0$ or $N/T\rightarrow\infty$. We illustrate the IPP and its consequence to the structural parameter estimator in the example below, together with the issue of $N/T\rightarrow\kappa$.
Example 3 (heterogeneous mean model)
Following the same setting in Example 2 but let $\varepsilon_{it}$ be normal with mean 00 and variance $\sigma_{0}>0$. In the current example, $\theta_{0}=(\mu_{0},\sigma_{0})^{\prime}$ is the structural parameter of interests. This example is a modified version of Neyman and Scott (1948) and is similar to Fernández-Val and Weidner (2016), except that the mean of $Y_{it}$ is $\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}$, conditional on $\phi^{0}$.
Define [^4] In this simple example, $\varepsilon_{it}$ is independent across both $i$ and $t$. Therefore, $\mathbb{E}$ is simply w.r.t. the normal density having mean $\mu^{0}+\alpha_{i}^{0}+\gamma_{t}^{0}$ and variance $\sigma^{0}$. $\widetilde{Y}_{it}:=Y_{it}-\mathbb{E}Y_{it}$. Following the same steps in Example 2, we can derive
$\widehat{\alpha}_{i}-\alpha_{i}^{0}$ $=\frac{1}{T}\sum_{t=1}^{T}\widetilde{Y}_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\qquad i=1,\ldots,N-1,$
$\widehat{\gamma}_{t}-\gamma_{t}^{0}$ $=\frac{1}{N}\sum_{i=1}^{N}\widetilde{Y}_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\qquad t=1,\ldots,T-1,$ (2.15)
where $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ can be obtained by maximizing $\mathbb{E}l(\mu,\psi)$ in the same way. Denote additionally
$\widehat{\alpha}_{N}:=-\sum_{i=1}^{N-1}\widehat{\alpha}_{i},\qquad\alpha_{N}^{0}:=-\sum_{i=1}^{N-1}\alpha_{i}^{0},\qquad\widehat{\gamma}_{T}:=-\sum_{t=1}^{T-1}\widehat{\gamma}_{t},\qquad\gamma_{T}^{0}:=-\sum_{t=1}^{T-1}\gamma_{t}^{0}.$
It follows that
$\widehat{\alpha}_{N}-\alpha_{N}^{0}=\frac{1}{T}\sum_{t=1}^{T}\widetilde{Y}_{Nt}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\qquad\widehat{\gamma}_{T}-\gamma_{T}^{0}=\frac{1}{N}\sum_{i=1}^{N}\widetilde{Y}_{iT}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it}.$ (2.16)
Now, for $\psi^{0}:=({\alpha}_{1}^{0},\ldots,{\alpha}_{N-1}^{0},{\gamma}_{1}^{0},\ldots,{\gamma}_{T-1}^{0})^{\prime}$, $\widehat{\psi}:=(\widehat{\alpha}_{1},\ldots,\widehat{\alpha}_{N-1},\widehat{\gamma}_{1},\ldots,\widehat{\gamma}_{T-1})^{\prime}$, and $Y_{it}(\mu):=Y_{it}-\mu-\alpha_{i}^{0}-\gamma_{t}^{0}$, we can calculate
$l(\theta,\widehat{\psi})-l(\theta,\psi^{0})$ $=\frac{1}{\sigma}\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\left(Y_{it}(\mu)-\frac{1}{2}(\widehat{\alpha}_{i}-\alpha_{i}^{0})\right)(\widehat{\alpha}_{i}-\alpha_{i}^{0})$
$\quad\quad+\frac{1}{\sigma}\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\left(Y_{it}(\mu)-\frac{1}{2}(\widehat{\gamma}_{t}-\gamma_{t}^{0})\right)(\widehat{\gamma}_{t}-\gamma_{t}^{0}),$ (2.17)
where $\theta=(\mu,\sigma)^{\prime}$ with $\sigma>0$. Substituting (2.15) and (2.16) into (2.17), we obtain, after some algebra,
$l(\theta,\widehat{\psi})-l(\theta,\psi^{0})$ $=\frac{1}{2\sigma}\frac{1}{NT^{2}}\sum_{i=1}^{N}\sum_{t=1}^{T}\sum_{s=1}^{T}\widetilde{Y}_{it}\widetilde{Y}_{is}+\frac{1}{2\sigma}\frac{1}{N^{2}T}\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it}\widetilde{Y}_{jt}$
$\quad\quad-\frac{2}{\sigma}\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\left(\mu-\mu^{0}\right)\widetilde{Y}_{it}+O_{p}(1/(NT)).$
Here, $\mathbb{E}\widetilde{Y}_{it}=0$ by definition and $\mathbb{E}\widetilde{Y}_{it}\widetilde{Y}_{js}=0$, for $i\neq j$ or $t\neq s$, by independence. Therefore, calculating the expectation, we arrive at